# BemaGANv2: A Tutorial and Comparative Survey of GAN-based Vocoders for Long-Term Audio Generation

### v1

@@ 4 Experiments and Results
In this section, we conduct a comprehensive evaluation of various discriminator configurations and their impact on audio reconstruction quality. We place particular emphasis on the architectural combinations used in BemaGANv2, while also referencing models such as HiFi-GAN, BigVGAN, and our previous BemaGAN. The goal is not only to compare model performance, but also to offer practical insights for vocoder design choices in short-term and long-term audio synthesis tasks. For short-term evaluation, we used audio clips of approximately 20 seconds, including a diverse range of content such as human speech, drum and percussion sounds, short musical phrases, and car engine noise. For long-term evaluation, we used 90-second segments of full music tracks spanning various genres. To ensure a fair comparison of discriminator configurations, all models except HiFi-GAN employed a generator architecture based on AMP blocks with embedded Snake activation functions.
@@ 4.1 Evaluation Metrics
  • Pearson Correlation Coefficient (PLCC): Measures the degree of linear correlation in frequency domain characteristics[15].
@@
In our implementation, we adopt the standard SPTK-based configuration and compute MCD using fastDTW after extracting the mel-cepstrum from each frame. While the reported MCD values in our table may exceed 1 due to the evaluation on relatively long-term (30s, 90s) audio and due to practical constraints, such as limited computing resources, short training epochs, and the need to evaluate over six different model configurations. However, when we test under typical conditions used in prior work—namely short-form utterances of around 10 seconds and sufficiently trained models—we are able to achieve MCD values well below 1.0, which we have confirmed empirically. This validates the effectiveness and correctness of our MCD implementation.
@@ 4.2 Objective Evaluation Results
Models FAD $\downarrow$ SSIM $\uparrow$ PLCC $\sim 1$ MCD $\downarrow$
BemaGANv2 0.911 0.78 0.949 1.652
BemaGAN 3.194 0.70 0.861 3.18
BigVGAN 1.906 0.71 0.906 2.56
HiFi-GAN 6.06 0.73 0.792 3.22
MSD + MED 2.337 0.68 0.879 2.964
MSD + MRD 3.327 0.66 0.841 3.354
MED + MPD + MRD 2.146 0.71 0.851 3.136
@@
Models FAD $\downarrow$ SSIM $\uparrow$ PLCC $\sim 1$ MCD $\downarrow$
BemaGANv2 2.685 0.39 0.945 1.8
BemaGAN 6.612 0.43 0.866 2.841
BigVGAN 3.58 0.35 0.908 2.28
HiFi-GAN 30.883 0.53 0.702 7.75
MSD + MED 6.66 0.35 0.810 6.92
MSD + MRD 18.27 0.39 0.776 7.52
MED + MPD + MRD 5.969 0.40 0.878 2.7557
While our earlier experiments indicated that HiFi-GAN performed better in short-duration tasks, additional training with extended epochs has revealed a different trend[3]. BemaGANv2, when trained for a longer duration, not only maintains its superior performance in long-term audio generation but also clearly outperforms other models in short-term tasks.
We hypothesize that this improvement is linked to the gradient behavior of the Snake activation function. Unlike Leaky ReLU, which maintains a non-zero gradient across its domain, Snake contains regions where the derivative approaches zero, potentially slowing convergence in early training stages. However, with sufficient training, the model leverages the periodic inductive bias of Snake to capture harmonic and transient features more effectively.
A detailed mathematical explanation of this behavior, including derivative analysis, is provided in  A.
@@ 4.3 Subjective Evaluation Results
Table 4 and Table 5 report the results of human evaluations, which are broadly consistent with the objective measurements. These results provide further evidence that architectural decisions—such as the choice of activation functions and discriminator types—impact the perceptual quality across different audio durations.
@@ 4.5 Result
The results clearly demonstrate that BemaGANv2 achieves superior performance in both short- and long-term audio tasks, outperforming all other models in objective and subjective evaluations. This highlights the effectiveness of its generator design and discriminator configuration, particularly in preserving both high-frequency details and long-term periodic structures.
Although BemaGANv2 shows relatively low SSIM scores, this can be attributed to its behavior in reconstructing amplitude levels. SSIM is computed by comparing Mel-spectrogram images across three RGB channels, where the color intensity primarily reflects the sound loudness. The lower SSIM indicates that BemaGANv2 tends to deviate from the exact decibel scale of the original audio. However, subjective metrics such as MOS and SMOS consistently rank BemaGANv2 as the top performer. This suggests that human listeners are more sensitive to the perceptual fidelity of the audio, such as timbral and structural quality, than to precise restoration of loudness. In support of this, BemaGANv2 also achieves the best scores in FAD and MCD, which are directly related to perceptual realism and spectral coherence.
Interestingly, although the MED+MPD+MRD combination yielded decent objective metrics — especially in long-term evaluation — it showed significantly lower subjective scores. Through an analysis of training and validation loss curves, we confirmed that the issue is not due to overfitting. Instead, we attribute this degradation in perceptual quality to mode collapse[17], likely caused by adversarial instability when employing an excessive number of discriminators. Following this observation, we did not pursue further experiments involving combinations of three or more discriminators, as the risk of mode collapse outweighed potential performance gains.
These findings emphasize the importance of balancing feedback diversity and adversarial stability in discriminator design. For further experimental details, including loss curves and configuration analysis, please refer to  B.
@@ 5 Discussion
  • Model architecture — no differences except the activation function.
The only significant difference between HiFi-GAN and the other models lies in the activation function: HiFi-GAN utilizes LeakyReLU[2], while others incorporate the Snake function. Given that Snake is known to better extrapolate periodic structures, especially in long audio sequences[5], we conclude that the root cause of the abnormal doubling in waveform duration stems from HiFi-GAN’s inability to generalize beyond the training distribution of short-length signals.
This result underscores the critical role of periodic inductive bias in vocoders tasked with long-term synthesis. Snake-based models consistently exhibit better stability and fidelity.
@@ 6 Conclusion
In this study, we presented BemaGANv2, a vocoder designed to enhance both the fidelity and temporal coherence of waveform reconstruction. Through theoretical motivations and extensive empirical evaluations, we demonstrated that BemaGANv2 outperforms existing GAN-based vocoders such as HiFi-GAN and BigVGAN, particularly in long-duration audio tasks. The model’s generator architecture incorporating AMP blocks and the proposed Multi-Envelope Discriminator (MED) proved effective in capturing both local details and long-term periodicity.
Our results showed that BemaGANv2 achieves superior performance in both objective and subjective evaluation metrics in short- and long-term audio. Notably, it exhibited perceptual advantages even in scenarios where traditional metrics such as SSIM were lower, highlighting its ability to preserve the auditory quality that human listeners prioritize.
These findings collectively suggest that BemaGANv2 is a highly suitable vocoder for Text-to-Music (TTM) and Text-to-Audio (TTA) models, especially diffusion based generation pipelines. As a next step, we plan to integrate BemaGANv2 into multimodal generative AI systems, enabling more expressive and temporally coherent audio generation conditioned on various modalities. We believe this direction will further bridge the gap between machine-generated and human-perceived audio realism.
@@ Appendix A Analysis of the Activation Function
Our earlier experiments suggested that HiFi-GAN performs best on short-duration audio[3], especially under limited training epochs. However, in the extended study presented in this paper, we provided sufficient training epochs, and the results indicate that BemaGANv2 outperforms all baselines in both short- and long-term waveform reconstruction tasks. In this section, we provide a mathematical interpretation to explain this behavior, particularly in relation to the activation functions used.
@@ C.3 Mel-to-Audio Consistency Verification
[…]

### v3

@@ 4 Experiments and Results
In this section, we conduct a comprehensive evaluation of various discriminator configurations and their impact on audio reconstruction quality. We place particular emphasis on the architectural combinations used in BemaGANv2, while also referencing models such as HiFi-GAN, BigVGAN, and our previous BemaGAN. The goal is not only to compare model performance, but also to offer practical insights for vocoder design choices in short-term and long-term audio synthesis tasks. This section is structured as follows: we first outline the evaluation metrics used, then present the objective and subjective results, followed by a qualitative comparison of Mel-spectrogram structures, and finally summarize our key findings. For short-term evaluation, we used audio clips of approximately 20 seconds, including a diverse range of content such as human speech, drum and percussion sounds, short musical phrases, and car engine noise. For long-term evaluation, we used 90-second segments of full music tracks spanning various genres. It is important to note that both the short-term and long-term test sets were intentionally constructed to evaluate out-of-distribution (OOD) generalization, as all test audio—including environmental sounds, musical instruments, and full music tracks—differs substantially from the single-speaker speech in the LJSpeech training corpus. This represents a significant domain shift and provides a rigorous test of each model’s extrapolation capabilities beyond the training distribution. To ensure a fair comparison of discriminator configurations, all models except the original HiFi-GAN employed a generator architecture based on AMP blocks with embedded Snake activation functions. All models in the comparative study were trained for 500 epochs under identical conditions (detailed in Appendix D.7). Following the established findings from HiFi-GAN [2] and BigVGAN [3], which demonstrated that multi-discriminator configurations significantly outperform single-discriminator setups, our primary experiments focus on evaluating novel combinations of discriminators (MED+MRD, MSD+MED, etc.).
In addition, to address the need for empirical validation of individual component contributions, we conducted two targeted ablation experiments. First, we evaluated a MED-only configuration, where the AMP+Snake generator is paired with MED as the sole discriminator. This isolates MED’s independent contribution to audio quality, particularly in temporal envelope modeling. Second, we applied the HiFi-GAN discriminator combination (MPD+MSD) to the same AMP+Snake generator used across all other models. This configuration, denoted as HiFi-GAN w/ AMP+Snake in our tables, strengthens variable control by holding the generator constant and allowing readers to directly compare the effect of different discriminator configurations. Note that a detailed ablation study on the Snake activation function was previously conducted by Lee et al. [3], where Snake was shown to significantly improve periodicity modeling compared to Leaky ReLU. Rather than duplicating this established result, the inclusion of the HiFi-GAN w/ AMP+Snake experiment indirectly illustrates the generator-side contribution by contrasting it with the original HiFi-GAN, which uses the Leaky ReLU-based ResBlock generator.
It is important to emphasize that the goal of our comparative experiments is not to identify which single discriminator performs best in isolation, but rather to demonstrate that how discriminators are combined matters as much as—if not more than—the choice of individual discriminators. The ablation experiments complement this primary objective by providing empirical evidence for the individual roles of key components.
@@ 4.1 Evaluation Metrics
  • Pearson Correlation Coefficient (PCC): Measures the degree of linear correlation in frequency domain characteristics[17].
@@
  • Multi-Resolution STFT Loss (M-STFT): Measures spectral reconstruction quality across multiple STFT resolutions [42].
  • Periodicity: Evaluates the preservation of periodic structures in the generated waveform by measuring pitch periodicity estimation error [43].
Among these metrics, M-STFT and Periodicity are adopted following the evaluation protocol established by BigVGAN [3], providing a standardized framework for spectral reconstruction quality and harmonic structure preservation that complements the existing perceptual and distributional metrics.
We note that certain metrics used in BigVGAN—specifically PESQ (Perceptual Evaluation of Speech Quality, ITU-T P.862) and V/UV F1 (Voiced/Unvoiced F1 score)—were not included in our evaluation. Both metrics are designed exclusively for speech signals: PESQ models the human auditory system’s sensitivity to speech distortions, while V/UV F1 measures classification accuracy of voiced and unvoiced segments. Since our test sets include musical content, environmental sounds, and other non-speech audio, these speech-specific metrics would not yield reliable or interpretable results across the entire evaluation set.
Regarding FAD, while it remains widely used for evaluating generative audio models, recent studies have identified several limitations. Gui et al. [41] demonstrated that FAD scores are susceptible to sample size bias, sensitive to the choice of audio embedding model, and influenced by the quality and composition of the reference set. Their findings further showed that commonly used configurations—such as VGGish embeddings with the MusicCaps reference set—may not correlate well with subjective quality judgments. In light of these concerns, we retain FAD as a supplementary distributional measure for completeness and comparability with prior work, but avoid overreliance on FAD as primary evidence of perceptual quality.
In our implementation, we adopt the standard SPTK-based configuration and compute MCD using fastDTW after extracting the mel-cepstrum from each frame. It is important to note that while typical MCD values for high-quality speech synthesis are often below 1.0, the reported MCD values in our tables, may exceed 1. This is attributed to practical constraints, such as limited computing resources, short training epochs because of the need to evaluate across six different model configurations. However, we empirically confirmed that when tested under typical conditions used in prior work—namely short-form utterances of around 10 seconds and with sufficiently trained models—our implementation yields MCD values well below 1.0. This validates the effectiveness and correctness of our MCD implementation for comparative analysis across the evaluated models.
SSIM, originally designed for image comparison, has been successfully adapted for audio evaluation by measuring structural similarity between mel-spectrogram representations [37, 38]. It effectively captures spectral-temporal coherence and has been validated in recent vocoder benchmarks.
Similarly, we use Pearson Correlation Coefficient (PCC) which measures correlation between ground-truth and reconstructed mel-spectrograms. PCC has been widely adopted in audio reconstruction tasks [39, 40] and correlates well with perceptual similarity.
@@ 4.2 Objective Evaluation Results
Models FAD $\downarrow$ SSIM $\uparrow$ PCC $\sim 1$ MCD $\downarrow$ M-STFT $\downarrow$ Periodicity $\downarrow$ BemaGANv2 (MED + MRD) 0.911 0.85 0.949 1.652 1.262 0.161    MED only 1.116 0.82 0.94 1.887 1.453 0.2208 BemaGAN (MPD + MED) 3.194 0.68 0.861 3.18 1.893 0.1883 BigVGAN (MPD + MRD) 1.906 0.79 0.906 2.56 1.519 0.1716 HiFi-GAN (MPD + MSD) 6.06 0.73 0.792 3.22 3.256 0.1883    w/ AMP + Snake 1.5363 0.78 0.888 2.553 1.527 0.1543 MSD + MED 2.337 0.67 0.879 2.964 1.803 0.1735 MSD + MRD 3.327 0.63 0.841 3.354 1.948 0.1802 MED + MPD + MRD 2.146 0.71 0.851 3.136 1.771 0.1692
@@
Models FAD $\downarrow$ SSIM $\uparrow$ PCC $\sim 1$ MCD $\downarrow$ M-STFT $\downarrow$ Periodicity $\downarrow$ BemaGANv2 (MED + MRD) 2.681 0.78 0.945 1.8 1.5141 0.1235    MED only 2.204 0.75 0.945 1.966 1.638 0.1361 BemaGAN (MPD + MED) 6.612 0.63 0.866 2.841 1.827 0.1657 BigVGAN (MPD + MRD) 3.58 0.71 0.908 2.28 1.613 0.1504 HiFi-GAN (MPD + MSD) 30.883 0.08 0.702 7.75 3.115 0.2178    w/ AMP + Snake 4.274 0.69 0.885 2.392 1.622 0.1483 MSD + MED 6.66 0.63 0.810 2.92 1.7301 0.1922 MSD + MRD 7.221 0.60 0.776 3.345 1.75 0.1793 MED + MPD + MRD 5.969 0.66 0.878 2.7557 1.6801 0.1762
As presented in Table 2 and Table 3, BemaGANv2 (MED+MRD) consistently achieves superior performance across most objective metrics for both short- and long-term audio generation. BemaGANv2 outperforms all baselines in FAD, SSIM, PCC, MCD, and M-STFT for short-term audio, and achieves the best scores in SSIM, PCC, MCD, M-STFT, and Periodicity for long-term audio.
HiFi-GAN with its original generator (Leaky ReLU + ResBlock) shows significantly degraded performance in long-term audio across all metrics, which aligns with the anomalous waveform length doubling observed in our extended analysis (Section 5). In contrast, when the same HiFi-GAN discriminator combination (MPD+MSD) is applied to the AMP+Snake generator (HiFi-GAN w/ AMP+Snake), performance improves substantially—e.g., long-term FAD drops from 30.883 to 4.274, and MCD from 7.75 to 2.392. This result provides strong evidence that the generator architecture, particularly the Snake activation and anti-aliasing mechanism, plays a critical role in long-term audio quality, consistent with the findings of Lee et al. [3]. Notably, the duration doubling anomaly previously observed in HiFi-GAN was not present in this configuration, further supporting the generator-side origin of the issue.
Ablation: MED-only configuration. To isolate the contribution of MED, we evaluated a configuration using MED as the sole discriminator with the same AMP+Snake generator. The MED-only model achieves competitive results: in short-term evaluation, it attains the second-lowest M-STFT (1.453) and strong PCC (0.94), and in long-term evaluation, it achieves the lowest FAD (2.204) among all models. These results confirm that MED independently contributes meaningful supervisory signals for audio quality, particularly in distributional fidelity and temporal envelope modeling. However, compared to the full BemaGANv2 (MED+MRD), the MED-only configuration shows lower performance in SSIM (0.82 vs. 0.85 for short-term; 0.75 vs. 0.78 for long-term), M-STFT (1.453 vs. 1.126 for short-term), and Periodicity (0.2208 vs. 0.161 for short-term; 0.1361 vs. 0.1235 for long-term). This demonstrates that MRD provides complementary spectral supervision that enhances structural consistency and harmonic preservation—capabilities that MED’s temporal envelope analysis alone does not fully cover. The synergy between MED and MRD is thus empirically validated: MED excels at temporal envelope and distributional fidelity, while MRD strengthens spectral sharpness accuracy.
While our earlier experiments indicated that HiFi-GAN performed better in short-term tasks[6], additional training with extended epochs has revealed a different trend. BemaGANv2, when trained for a longer duration, not only maintains its superior performance in long-term audio generation but also clearly outperforms other models in short-term tasks. This trend may be attributed to the gradient properties of the Snake activation function. Unlike Leaky ReLU, which maintains a non-zero gradient across its domain, Snake contains regions where the derivative approaches zero, potentially slowing convergence in early training stages. However, with sufficient training, the model leverages the periodic inductive bias of Snake to capture harmonic and transient features more effectively. A detailed mathematical explanation of this behavior, including derivative analysis and its implications for convergence, is provided in A.
To verify the robustness of these findings, we additionally trained BemaGANv2 with four different random seeds and report the mean and standard deviation of all objective metrics. The results, presented in Appendix E, confirm that BemaGANv2 exhibits low variability across seeds, with most metrics showing small standard deviations, indicating that the reported performance advantages are reproducible and not dependent on a particular random initialization.
@@ 4.3 Subjective Evaluation Results
Table 4 and Table 5 report the results of human evaluations, which are broadly consistent with the objective measurements. These results provide further evidence that architectural decisions—such as the choice of activation functions and discriminator types—impact the perceptual quality across different audio durations. The MOS and SMOS evaluations were conducted with 137 general participants recruited via crowdsourcing. Each audio sample was rated by at least 15 listeners. All reported scores include 95% confidence intervals(CI) as shown in Table 4 and Table 5. We used general listeners rather than expert listeners to ensure our evaluation reflects real-world user perception, which is the ultimate goal of high-fidelity audio generation systems.
@@ 4.5 Result
The results clearly demonstrate that BemaGANv2 achieves superior performance in both short- and long-term audio tasks, outperforming all other models across objective and subjective evaluations. This highlights the effectiveness of its generator design and the MED+MRD discriminator configuration, particularly in preserving both high-frequency details and long-term periodic structures. The addition of ablation configurations—MED only and HiFi-GAN with the AMP+Snake generator—further confirms that the MED+MRD combination provides the most balanced supervisory signal, and that the generator architecture plays a critical role in long-term generation stability.
Interestingly, although the MED+MPD+MRD combination yielded competitive objective metrics—especially in long-term evaluation—it showed significantly lower subjective scores. Through an analysis of training and validation loss curves (Appendix B), we confirmed that this is not due to overfitting. Instead, we attribute the degradation in perceptual quality to mode collapse, likely caused by adversarial instability when employing an excessive number of discriminators. This observation, combined with the ablation results showing that even the MED-only configuration achieves reasonable performance, suggests that discriminator design should prioritize complementary coverage of perceptual dimensions over sheer quantity. These findings highlight a crucial practical insight: the importance of carefully balancing feedback diversity and adversarial stability in discriminator design, cautioning against simply adding more discriminators without considering their synergistic effects.
For further experimental details, including loss curves and configuration analysis, as well as complete implementation specifications—such as hardware settings, all hyperparameters, preprocessing pipelines, and training configurations—please refer to B and D. All code, configuration files, and pretrained models are publicly available at the link mentioned in the abstract.
@@ 5 Discussion
  • Model architecture — no differences except the generator architecture.
Among the models evaluated in our comparative study, HiFi-GAN is the only one that uses the Leaky ReLU-based ResBlock generator, while all other models employ the AMP+Snake generator architecture. This observation initially suggested the activation function as a potential factor in the anomaly. To investigate this further, we conducted an additional experiment in which HiFi-GAN’s discriminator combination (MPD+MSD) was applied to the same AMP+Snake generator used across all other models. Notably, when using this configuration (HiFi-GAN w/ AMP+Snake), the waveform duration doubling anomaly was not observed, and long-term audio quality improved substantially. This result provides strong evidence suggesting that the generator architecture—particularly the choice of activation function and the presence of anti-aliasing—plays a significant role in the observed anomaly, rather than the discriminator configuration.
These findings underscore the importance of periodic inductive bias in vocoders tasked with long-term synthesis. The Snake activation function, which provides learnable periodicity through its oscillatory formulation [4], appears to enable more stable extrapolation of periodic structures beyond the training distribution of short-length signals. However, we note that the AMP+Snake generator differs from HiFi-GAN’s original generator in multiple aspects—including both the activation function and the anti-aliasing mechanism—and therefore, while these results substantially narrow the source of the anomaly to the generator side, a fully isolated single-variable experiment (e.g., changing only the activation function while keeping all other architectural components identical) would be needed to establish definitive causality. Nevertheless, in all of our experiments, Snake-based generators consistently exhibited better stability and fidelity in long-term audio generation.
Beyond the HiFi-GAN anomaly discussed above, we also observed performance limitations of BemaGANv2 in certain audio conditions. Specifically, BemaGANv2 achieves high reconstruction quality on relatively simple audio sources such as solo instruments and single vocal tracks, but exhibits noticeable quality degradation when processing complex polyphonic audio with multiple simultaneously playing instruments (e.g., full band arrangements, orchestral passages). We attribute this to the increased harmonic overlap and spectral density in such signals, which pose challenges for the envelope-based discrimination provided by MED, combined with the fact that the model was trained on LJSpeech, which consists exclusively of single-speaker speech. Addressing this limitation through training on more diverse and polyphonic datasets represents an important direction for future work.
@@ 6 Conclusion
In this study, we presented BemaGANv2, a vocoder designed to enhance both the fidelity and temporal coherence of waveform reconstruction. Motivated by theoretical considerations and validated through extensive empirical evaluation, BemaGANv2 was shown to outperform existing GAN-based vocoders such as HiFi-GAN and BigVGAN, particularly in long-term audio generation. Its generator, based on AMP blocks, and the proposed Multi-Envelope Discriminator (MED), effectively capture both local signal detail and long-range periodic structure.
Our results demonstrate that BemaGANv2 achieves superior performance across both objective and subjective evaluation metrics, in both short- and long-form audio scenarios.
Key contributions of this work include:
  • A tutorial-style analysis of GAN-based vocoder evolution, from MelGAN to BigVGAN.
  • The design and integration of the Multi-Envelope Discriminator (MED) to enhance prosodic and envelope-level modeling.
  • A comparative evaluation of six discriminator configurations, shedding light on architectural trade-offs.
  • An empirical finding linking activation functions to inference stability in long-term generation tasks.
These findings suggest that BemaGANv2 is a promising vocoder for Text-to-Music (TTM) and Text-to-Audio (TTA) models, particularly in diffusion-based generation pipelines. As future work, we plan to integrate BemaGANv2 into multimodal generative AI systems to enable more expressive and temporally coherent audio synthesis. We anticipate that this direction will help bridge the gap between machine-generated and human-perceived audio realism.
A limitation of this study is that experiments were conducted primarily on the LJSpeech dataset, a single-speaker English corpus. While we attempted to scale to larger, multi-speaker datasets such as LibriTTS, the computational demands of our study—which requires training and evaluating multiple discriminator configurations under identical conditions—exceeded our available resources (a single NVIDIA A100 GPU via Google Colab, compared to the 8 NVIDIA V100 GPUs used in the BigVGAN study). Future work should include evaluation on multi-speaker and multi-language datasets such as LibriTTS and VCTK to further validate the generalizability of BemaGANv2 across diverse acoustic conditions.
@@ Appendix A Analysis of the Activation Function
Our earlier experiments suggested that HiFi-GAN performs best on short-term audio[6], especially under limited training epochs. However, in the extended study presented in this paper, we provided sufficient training epochs, and the results indicate that BemaGANv2 outperforms all baselines in both short- and long-term waveform reconstruction tasks. In this section, we provide a mathematical interpretation to explain this behavior, particularly in relation to the activation functions used.
@@ C.3 Mel-to-Audio Consistency Verification
  • $\text{mel\_length}=\text{waveform\_length}/\text{hop\_size}$